### v2

# ![[Uncaptioned image]](extracted/6365550/Fig/teaser.png) Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution

Abstract.
Chain-of-Thought (CoT) prompting enhances the reasoning of large language models (LLMs) by decomposing problems into sequential steps, mimicking human logic and reducing errors. However, complex tasks with vast solution spaces and vague constraints often exceed the capacity of a single reasoning chain. Inspired by Minimal Free Resolution (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a novel framework that extends CoT by introducing auxiliary, interrelated reasoning paths. SoT captures deeper logical dependencies, enabling more robust and structured problem-solving. MFR decomposes a module into a sequence of free modules with minimal rank, providing a structured analytical approach to complex systems. This method introduces the concepts of ”Module”, ”Betti numbers”,”Freeness”, ”Mapping”, ”Exactness” and ”Minimality”, enabling the systematic decomposition of the original complex problem into logically complete minimal subproblems while preserving key problem features and reducing reasoning length. We tested SoT across diverse datasets (e.g., GSM8K, MATH) and models (e.g., GPT-4o-mini, Qwen2.5), achieving inference accuracy that matches or surpasses mainstream CoTs standards. Additionally, by aligning the sampling process with algebraic constraints, our approach enhances the scalability of inference time in LLMs, ensuring both transparent reasoning and high performance. Our code will be publicly available at https://github.com/dlMARiA/Syzygy-of-thoughts
1. Introduction
The rapid advancements in natural language processing (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2018; Brown et al., 2020) have demonstrated outstanding reasoning capabilities across various tasks (Zhao et al., 2023b; Min et al., 2023). In recent years, the Chain of Thought (CoT) prompting method (Wei et al., 2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems (Chen et al., 2025; Qu et al., 2025; Sui et al., 2025; Wei et al., 2025; Xia et al., 2024; Liu et al., 2025). The core advantage of the CoT framework lies in its ability to decompose problems into a series of explicit and easily interpretable intermediate steps, thereby improving the transparency and explainability of the reasoning process. This explicit stepwise reasoning not only facilitates in-depth analysis and debugging of model outputs but also provides a clear logical pathway for solving complex problems.
However, despite its success in many tasks, CoT exhibits inherent limitations when dealing with high-dimensional, nonlinear, and abstract logical problems (Wang et al., 2022a; Shi et al., 2023; Jin et al., 2024). First, traditional CoT methods often rely on fixed or heuristic decomposition strategies, which struggle to capture essential details and intrinsic connections in multidimensional problems (Wang et al., 2022a). Second, due to the lack of systematic structural planning in intermediate steps, the reasoning process tends to include redundant computations and repetitive paths, leading to increased computational resource consumption and inference latency (Jin et al., 2024). Although existing improvements have alleviated these issues to some extent, introducing techniques such as multi-path exploration, voting mechanisms, and hierarchical decomposition strategies to enhance robustness (Yao et al., 2023a; Besta et al., 2024; Wang et al., 2022b; Zhou et al., 2022; Yao et al., 2023b), however, they have yet to fully address logical incompleteness and the inadequate decomposition of high-dimensional information. As a result, when handling highly complex reasoning tasks, model robustness and accuracy still face significant challenges.
To address these challenges, we propose a novel reasoning framework—Syzygy of Thoughts (SoT). This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra (Eisenbud, 2013; Rossi and Sharifan, 2009; Botbol et al., 2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences. The incorporation of MFR theory ensures a more rigorous and efficient problem decomposition process while preserving the core characteristics of the problem, effectively reducing redundant computations and logical inconsistencies. Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as ”module”, ”Betti numbers”, ”freeness”, ”mapping”, ”exactness” and ”minimality” to deeply explore and systematically express the intrinsic structure of complex problems. This approach provides a novel theoretical perspective for analyzing high-dimensional and multivariable problems.
To validate the effectiveness of SoT, we conducted extensive experiments across diverse datasets, including GSM8K, MATH, and others, spanning mathematical reasoning and high-dimensional problem domains. These datasets, with their complex structures and multifaceted features, provide a robust testbed for evaluating reasoning frameworks in real-world scenarios. Experimental results show that SoT achieves inference accuracy that matches or surpasses mainstream CoTs standards, demonstrating its superior performance and practical potential.
Finally, the SoT framework exhibits multiple advantages, including an intuitive and transparent reasoning path and improved interpretability of intermediate information. The main contributions of this work are as follows:
Pioneering the integration of MFR into LLM reasoning, enabling structured decomposition of complex problems.
Developing a novel framework that builds on CoT’s strengths while substantially improving reasoning clarity and structural efficiency.
Providing theoretical insights and empirical evidence that SoT significantly enhances performance in high-dimensional reasoning tasks.
2. Related Work
2.1. Advanced Reasoning Techniques in LLMs
The reasoning capabilities of LLMs have advanced significantly with breakthroughs in prompting paradigms, notably CoT prompting proposed by Wei et al.(Wei et al., 2022). CoT decomposes complex problems into stepwise solutions, mimicking human reasoning, yet struggles with multidimensional or nonlinear tasks like abstract logic and advanced mathematics. To address these limitations, variants such as Zero-shot-CoT(Kojima et al., 2022), Self-Consistency CoT (Wang et al., 2022b), and Auto-CoT (Zhang et al., 2022) have emerged, alongside verification-based approaches like VerifyCoT (Zhao et al., 2023a) and CoF-CoT (Nguyen et al., 2023), yielding measurable improvements.
Structured prompting strategies further enhance reasoning. Least-to-Most Prompting (LtM)(Zhou et al., 2022) breaks tasks into manageable subproblems, reducing single-step errors. Programmatic methods, including Program of Thought (PoT)(Chen et al., 2022), Chain of Code (CoC)(Li et al., 2023), and Buffer of Thought (BoT)(Yang et al., 2025), integrate discrete variables and procedural steps. Building on these, Algorithm of Thought (AoT) (Sel et al., 2023) iteratively synthesizes and refines solutions, boosting token efficiency.
Tree- and graph-based paradigms offer alternative exploration. Tree-of-Thought (ToT)(Yao et al., 2023a) employs a hierarchical structure to explore multiple decision paths, though exponential growth limits deep efficiency. Graph-of-Thought (GoT)(Besta et al., 2024) extends this with flexible graph configurations, enabling path recalibration and information aggregation. Meanwhile, Skeleton-of-Thought (SoT)(Ning et al., 2023) generates high-level outlines before detailing, balancing speed and quality, while recursive methods like Self-Refine(Madaan et al., 2023) and Step-Back Prompting (Zheng et al., 2023) refine outputs iteratively for higher accuracy.
2.2. Minimal Free Resolution
MFR is a core tool in homological algebra (Fieldsteel and Nagel, 2021) and algebraic geometry (Botbol et al., 2021), used to analyze the algebraic structure of modules (Eisenbud, 2013; Rossi and Sharifan, 2009), revealing properties such as rank, symmetry, and relationships (Chen, 2011). In computational algebraic geometry, MFR analyzes singularities and invariants of algebraic varieties through syzygies (Evans and Griffith, 1981), optimizing the complexity and efficiency of Gröbner basis computations (Eisenbud et al., 2001; Stewart, 1993; Weispfenning, 1992; Capani et al., 1997).
In Topological Data Analysis (TDA), MFR accelerates persistent homology computations (Wasserman, 2018), enhancing the stability of topological analysis for high-dimensional data. In physics and bioinformatics, MFR is used to analyze Calabi-Yau manifold singularities (Polchinski, 1994), gauge field structures (Hitchin, 2003), and the modular properties of gene regulatory networks (Li et al., 2011).
The advantages of MFR in optimizing computational complexity and revealing algebraic structures provide a new perspective for its integration into the study of CoT in LLM. In CoT, structured reasoning steps are required, but often face challenges of inefficiency and redundancy. We innovatively introduce MFR into CoT by decomposing the symbolic dependencies between reasoning steps, eliminating redundant calculations, optimizing intermediate processes, and enhancing transparency. This approach opens up a more efficient and structured technical pathway for LLM applications in solving complex problems.
3. Method
3.1. Preliminaries
In algebraic geometry and homological algebra, minimal free resolution is an essential tool for studying the structure and properties of modules, especially when dealing with finitely generated modules over a ring.
Key Properties: In algebraic geometry and homological algebra, a ”syzygy” refers to the relations among the relations that generate a module. A syzygy is a relation among the generators. For example, if a module is defined by generators , then a syzygy is a relation of the form:
where . These relations themselves can form a module, which is called the ”first syzygy module.” A module over a ring is an algebraic structure similar to a vector space, but the scalars come from instead of a field. A module is called free if it has a basis, i.e., for some dimension (finite or infinite). Free modules are analogous to free vector spaces in linear algebra. A module is finitely generated if there exists a finite set such that . A MFR of a finitely generated module over is an exact sequence:
Where each is free (), the matrices of the maps contain no units, and the resolution is ”minimal,” meaning the number of free generators is the smallest possible. This resolution provides a standardized way to describe the structure of , reflecting both its complexity and its homological properties. A MFR is characterized by its minimality, where each uses the smallest number of generators, and the matrices in the maps contain no units. It is unique up to isomorphism, meaning any two MFRs of are naturally isomorphic. The -th Betti number of , defined as , counts the number of free generators in and reflects the complexity of the resolution.
3.2. Minimal Free Resolution enhanced CoT
As shown in Figure 3, the SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs. Each reasoning step—such as problem interpretation, task decomposition, subgoal chaining, and final inference—is abstracted as generators or relations within modules, reinterpreting CoT as algebraic objects with generative structures and latent dependencies, rather than merely a flat sequence of symbols.
![[Uncaptioned image]](extracted/6365550/Fig/modul.png)
Module: Initial Complex Problem
Complex problems often exhibit high-dimensional structures, intricate logical dependencies, and ambiguous constraints, rendering direct solutions intractable. In this initial module, we conceptualize the problem as a unified entity of interdependent substructures, necessitating systematic decomposition. The goal is not immediate resolution but to unravel latent complexity, identify core logical components, and establish a foundation for subsequent modular reasoning. This stage ensures that downstream processes operate within well-defined boundaries, setting the stage for structured analysis.
![[Uncaptioned image]](extracted/6365550/Fig/betti.png)
Betti Numbers: Quantifying Complexity
To measure decomposition complexity, we introduce Betti numbers as a quantitative metric, where each number reflects the count of auxiliary conditions at a reasoning level. Higher Betti numbers indicate greater intricacy, signaling potential optimization opportunities. We leverage LLMs to minimize these numbers by regenerating or filtering conditions, streamlining the decomposition process. Betti numbers thus guide reasoning efficiency, balancing structural complexity with computational simplicity.
![[Uncaptioned image]](extracted/6365550/Fig/freeness.png)
Freeness
: Auxiliary Conditions
Freeness involves generating auxiliary conditions to simplify problems and clarify logical relationships. By introducing intermediate variables, hypotheses, or formalized constraints, complex problems are partitioned into smaller, independent subproblems. This reduces dimensionality and provides clear entry points for resolution, enhancing the tractability of subsequent steps while maintaining logical coherence.
![[Uncaptioned image]](extracted/6365550/Fig/mapping.png)
Mapping: Solving Strategies
Effective problem-solving requires systematic strategies that map auxiliary conditions to actionable reasoning paths. These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises). LLMs iteratively solve subproblems along these paths, linking current conditions to new conclusions, ensuring continuity and interpretability throughout the reasoning chain.
![[Uncaptioned image]](extracted/6365550/Fig/exactness.png)
Exactness: Logical Completeness
Logical completeness is critical for rigor, preventing gaps or leaps in the reasoning chain. Each step must derive from explicit premises, with implicit assumptions formalized to avoid invalid conclusions. LLMs verify consistency across auxiliary conditions, subproblems, and the final solution, systematically reviewing the chain against original constraints. This ensures a coherent and reliable framework, minimizing errors from incomplete logic.
![[Uncaptioned image]](extracted/6365550/Fig/mini.png)
Minimality: Optimal Solutions
Minimality optimizes efficiency by deriving solutions with the fewest auxiliary conditions and simplest strategies, without sacrificing correctness. Redundant conditions are filtered, and reasoning paths are pruned of unnecessary steps. LLMs prioritize straightforward approaches over complex transformations, yielding concise, clear solutions. This reduces overall complexity, enhancing both computational efficiency and solution transparency.
After introducing the above six concepts, LLMs can be directly stimulated in a conversational manner. Figure 4 illustrates the conversational framework of these six concepts. First, complex reasoning problems are defined as a free module. Then, through the Betti number and auxiliary conditions of degrees of freedom, the complex problem is reduced and decomposed. By analyzing both the module and degrees of freedom simultaneously, the model can construct multiple mappings to solve the problem. Finally, minimality and accuracy are used to select the optimal solution.
3.3. Syzygy of Thoughts
The process of LLM reasoning for complex problems can be viewed as solving a complex reasoning problem through MFR. The entire process is formalized via a MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
Moreover, we define the intermediate space as a free module constructed from basic reasoning units:
where:
- is the reasoning space (LLM Reasoning Space), - are the basic units (Tokens) used in the reasoning construction, - Each is induced by the LLM, representing a single-step logical deduction in the CoT.
3.3.1. SoT Reasoning.
The Algorithm 1 and Figure 3 explain the entire process of SoT. The SoT framework begins with an initial complex problem , characterized by high-dimensional structure, intricate logical dependencies, and ambiguous constraint relationships. We treat as a unified representation of interdependent substructures, setting the stage for systematic decomposition. To analyze the inherent complexity of , we introduce Betti numbers as quantitative indicators of structural complexity. Specifically, the Betti number corresponds to the number of auxiliary conditions required for decomposition, with higher values suggesting deeper structural intricacy. These values guide the subsequent generation of auxiliary components and indicate potential optimization points.
We then introduce the notion of Freeness, which refers to the generation of auxiliary conditions that serve to simplify and structurally reorganize the problem. These conditions—denoted as —include intermediate hypotheses, latent constraints, or explicit sub-conclusions, and are generated via autoregressive reasoning using LLMs conditioned on previously established states. Each decomposes the original problem into finer-grained logical units, reducing its dimensional complexity and establishing solvable subspaces.
To resolve the decomposed problem, the framework constructs a set of mappings , each corresponding to a distinct reasoning path. These mappings are designed to satisfy two essential properties: directness, ensuring that each inference step avoids redundant operations, and logical soundness, ensuring that every conclusion is firmly grounded in its premises. Each mapping is generated using LLMs to traverse from auxiliary conditions to intermediate or final conclusions, thereby forming a continuous and interpretable reasoning chain.
To ensure logical completeness (Exactness) of the reasoning process, all steps within each reasoning path are verified for inferential closure. Implicit assumptions are formalized, and the consistency of auxiliary conditions, mappings, and final conclusions is evaluated. This validation step leverages the LLM’s reasoning and scoring capabilities via a scoring function . Finally, the optimal reasoning sequence is parsed using structured pattern-matching strategies (e.g., regular expressions, JSON parsing) to extract the final solution.
Method | Math Reasoning | Gene. | Multitask | Temporal | Log. | ||||
GSM8K | SVAMP | MultiArith | ASDiv | AQUA | MMLU | BBH | Date | CLUTRR | |
GPT-4o-mini | |||||||||
CoT | 85.1% | 84.4% | 99.2% | 97.0% | 65.0% | 63.1% | 66.3% | 51.8% | 65.9% |
CoT-SC (n=5) | 90.1% | 86.0% | 99.5% | 98.5% | 70.9% | 67.3% | 69.2% | 54.9% | 72.4% |
SoT (Ours) | 96.0% | 92.2% | 99.7% | 99.8% | 75.6% | 75.2% | 72.8% | 75.2% | 75.7% |
Qwen2.5-Coder-7B-Instruct | |||||||||
CoT | 77.2% | 82.4% | 92.3% | 92.0% | 60.6% | 55.1% | 47.1% | 31.0% | 20.1% |
CoT-SC (n=5) | 80.2% | 84.1% | 95.0% | 95.0% | 62.2% | 56.3% | 49.3% | 32.9% | 21.0% |
SoT (Ours) | 89.1% | 90.6% | 97.0% | 99.8% | 63.3% | 57.1% | 57.3% | 36.2% | 26.3% |
Qwen2.5-VL-72B-Instruct | |||||||||
CoT | 86.1% | 86.9% | 98.8% | 98.0% | 81.1% | 80.1% | 77.3% | 75.2% | 70.1% |
CoT-SC (n=5) | 89.1% | 88.2% | 99.3% | 98.4% | 83.9% | 82.9% | 79.0% | 78.0% | 75.0% |
SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
Gemma-3-27b-it | |||||||||
CoT | 83.1% | 85.9% | 91.9% | 98.5% | 80.3% | 70.8% | 70.7% | 76.9% | 65.3% |
CoT-SC (n=5) | 87.1% | 87.0% | 92.3% | 99.2% | 85.4% | 73.2% | 73.2% | 80.2% | 66.4% |
SoT (Ours) | 96.0% | 95.8% | 99.7% | 99.2% | 89.4% | 84.3% | 85.3% | 80.2% | 78.9% |
Gemma-3-12b-it | |||||||||
CoT | 83.2% | 79.0% | 90.4% | 97.7% | 68.9% | 68.1% | 64.6% | 77.7% | 49.0% |
CoT-SC (n=5) | 86.1% | 81.0% | 93.3% | 98.0% | 71.7% | 70.6% | 66.7% | 80.2% | 52.2% |
SoT (Ours) | 92.1% | 92.5% | 96.1% | 99.2% | 77.2% | 72.3% | 69.1% | 82.5% | 55.0% |
4. Experiments
To evaluate the performance of SoT, we designed a series of experiments, including applicability evaluation (Sec4.1), complex problem reasoning evaluation (Sec4.2), and ablation studies (Sec4.3, 4.4). First, the applicability evaluation tested SoT’s performance across LLMs of different scales, covering mathematical, planning, factual, and logical reasoning, thereby verifying its generalization and stability. Second, the complex problem reasoning evaluation focused on comparing SoT’s superiority in tackling complex tasks on the GSM8K and MATH datasets. Finally, the ablation studies analyzed key components, such as Betti numbers and temperature parameters, to examine their impact on the complexity, expressiveness, and robustness of the reasoning chain. These experiments comprehensively validate the theoretical and practical value of SoT.
4.1. Applicability Evaluation
The reasoning performance of chain-of-thought methods (such as CoT and CoT-SC) varies significantly across models of different sizes, especially on lightweight models. We aim to validate whether SoT can consistently enhance the reasoning capabilities of various LLMs and demonstrate its generalization and stability across models and task scenarios. For this purpose, we applied CoT, CoT-SC (n=5), and SoT to each model, with the default Betti number set to 7 and mapping set to 3, recording the accuracy across nine datasets.
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. We selected five different scales of LLMs, including GPT-4o-mini (lightweight), Qwen2.5-Coder-7B-Instruct (7B), Gemma-3-12b-it (12B), Gemma-3-27b-it (27B), and Qwen2.5-VL-72B-Instruct (72B), covering models from lightweight to large-scale. The baseline models were standard CoT and CoT-SC (n=5). The test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH (multi-task answering), Data (date prediction), and CLUTRR (logical reasoning), covering nine datasets in total.
Experimental Results: The results are shown in Table 1. SoT significantly outperforms the baselines across all models and datasets. For example, on GSM8K with GPT-4o-mini, SoT achieved 96.0%, improving by 10.9% over CoT (85.1%) and 5.9% over CoT-SC (90.1%). On the logical reasoning task CLUTRR, SoT achieved 75.7%, surpassing CoT (65.9%) and CoT-SC (72.4%). On Qwen2.5-VL-72B-Instruct, SoT achieved 89.4% on AQUA, outperforming CoT (81.1%) and CoT-SC (83.9%). For the lightweight model Qwen2.5-Coder-7B-Instruct, SoT improved GSM8K accuracy to 89.1%, 11.9% higher than CoT (77.2%), showing its superiority in resource-constrained scenarios. Furthermore, on Gemma-3-12b-it, SoT achieved 92.5% on SVAMP, close to the 95.8% achieved by Gemma-3-27b-it, indicating that its generalization ability is not limited by model scale. This result demonstrates the significant advantage of SoT’s structured design in improving reasoning performance and stability across models.
4.2. Complex Reasoning Evaluation
To compare the complex problem reasoning capabilities of SoT with mainstream methods, we evaluated SoT and compared it with extensive open-source data from various papers (Chen et al., 2025) . Specifically, we applied the SoT reasoning chain to GPT-4o-mini and tested its reasoning accuracy on the GSM8K and MATH datasets, while comparing it with other methods from the literature.
Experimental Configuration: The test model was GPT-4o-mini, with the Betti number set to 7 and mapping set to 3. The baselines included CoT and various popular chain-of-thought variants (such as MathPrompter, QuaSAR, MathDivide, etc.). The datasets used were GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
Method | Model | GSM8k | MATH |
No-CoT (Deng et al., 2024) | Mistral-7B | 38.0% | - |
ICoT-SI (Deng et al., 2024) | Mistral-7B | 51.0% | - |
- | RecurrentBlock-3.5B | 42.1% | - |
MathCoder-CL (Wang et al., 2023b) | Code-Llama-7B | 67.8% | 30.2% |
MAmmoTH (Yue et al., 2023) | Code-Llama-7B | 59.4% | - |
Brain (Chen et al., 2024a) | Code-Llama-7B | 74.0% | - |
SQ-VAE (Wang et al., 2023a) | Llama-2-7B | 40.0% | 7.0% |
Self-Rewarding (Chen et al., 2024b) | Llama-2-7B | 40.0% | 10.7% |
STaR (Zelikman et al., 2022) | Llama-2-7B | 58.2% | 16.0% |
ENVISIONS (Xu et al., 2024) | Llama-2-7B | 59.0% | 19.0% |
MetaMath (Yu et al., 2023b) | Llama-2-7B | 66.5% | - |
ToRA-Code (Gou et al., 2023) | Llama-2-7B | 72.6% | - |
OVM (Yu et al., 2023a) | Llama-2-7B | 73.7% | - |
- | Llama-3.1-8B | 56.7% | 20.3% |
- | Llama-3.1-70B | 85.5% | 41.4% |
- | Llama-3.1-405B | 89.0% | 53.8% |
- | NuminaMath-7B-CoT | 75.4% | 55.2% |
- | DeepSeek-Coder-7B | 77.4% | 44.4% |
- | Qwen2-7B | 79.9% | 44.2% |
- | Qwen2-Math-7B | 80.4% | 50.4% |
SIaM (Yu et al., 2024) | Qwen-2-Math-Base | 81.5% | 50.0% |
- | Internlm2-math-plus-7B | 84.0% | 54.4% |
OMI2 (Li et al., 2025) | Qwen2.5-Coder-7B | 84.1% | 72.3% |
CODEI/O++ (Li et al., 2025) | Qwen2.5-Coder-7B | 85.7% | 72.1% |
PyEdu (Li et al., 2025) | Qwen2.5-Coder-7B | 85.8% | 71.4% |
CODEI/O (Li et al., 2025) | Qwen2.5-Coder-7B | 86.4% | 71.9% |
OC-SFT-1 (Li et al., 2025) | Qwen2.5-Coder-7B | 86.7% | 70.9% |
WI (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.4% |
WI (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 87.0% | 71.1% |
OMI2 (Full) (Li et al., 2025) | Qwen2.5-Coder-7B | 88.5% | 73.2% |
- | DeepSeekMath-7B-RL | 88.2% | 51.7% |
CoMAT (Leang et al., 2024) | GPT-4 | 93.7% | - |
CoT (Ranaldi et al., 2025) | GPT-4 | 94.5% | - |
FCoT (Lyu et al., 2023) | GPT-4 | 95.0% | - |
MathPrompter (Imani et al., 2023) | GPT-4 | 95.6% | - |
QuaSAR (Radford et al., 2018) | GPT-4 | 96.5% | - |
MathDivide (Srivastava and Gandhi, 2024) | GPT-4 | 96.8% | - |
SoT (Ours) | GPT-4o-mini | 96.0% | 79.1% |
Experimental Results: As shown in Table 2, SoT achieved 96.0% on GSM8K and 79.1% on MATH on GPT-4o-mini. On GSM8K, SoT’s performance is close to the best result achieved by GPT-4 (e.g., MathDivide’s 96.8%) (Chen et al., 2025) and outperforms the best 7B model (e.g., OMI2 Full’s 88.5%) (Chen et al., 2025). On MATH, SoT’s 79.1% significantly outperforms the best mainstream 7B models (e.g., OMI2 Full’s 73.2%). This result indicates that SoT can achieve reasoning capabilities close to those of closed-source large models on lightweight models, significantly narrowing the performance gap between open-source models and GPT-4, validating its superiority in complex mathematical reasoning.
4.3. Betti Number Sensitivity
The SoT reasoning chain regulates topological constraints through the Betti number, and its value directly affects performance. To investigate the optimal configuration of the Betti number and its dynamic impact on structural expressiveness and regularization effects, we designed ablation experiments to reveal its key role in optimizing the reasoning process. For this, we systematically adjusted the Betti number and recorded accuracy changes under different mapping configurations. The goal of the experiment was to analyze the impact of the Betti number on the reasoning chain’s expressiveness and identify the performance saturation point. The results are presented in Figure 5, showing the trend of accuracy changes with varying Betti numbers, with particular attention to its non-monotonic properties and optimal configuration.
Experimental Configuration: The experiment was based on GPT-4o-mini and conducted on the GSM8K dataset. The Betti number was incremented from 1 to 10, with mapping set to 1, 2, and 3 to evaluate performance under different topological decomposition granularities.
Experimental Results: As shown in Figure 5, accuracy changed non-monotonically as the Betti number increased under mapping configurations of 1, 2, and 3. Specifically, when the Betti number started from 1, accuracy improved significantly, indicating that a small number of topological constraints effectively enhanced structural expressiveness. However, as the Betti number increased further, the performance gain gradually diminished and stabilized. The optimal saturation point was reached at 7, beyond which there was no noticeable improvement. The results reveal the dynamic regulatory effect of the Betti number in reasoning chain modeling, and that moderate topological constraints are crucial for optimizing SoT’s structured design, providing strong support for its theoretical foundation.
4.4. Stability Analysis
The temperature parameter influences the diversity of content generated by LLMs, which may challenge reasoning stability. We aimed to explore whether the SoT method could maintain its performance advantage under varying temperature conditions, thereby validating the adaptability of its structured design. To this end, we systematically evaluated SoT and CoT under different temperature conditions, focusing on the stability of accuracy and the relationship between reasoning diversity and consistency. The experiment involved adjusting the temperature step by step and observing the performance trends of the two methods under changing diversity.
Experimental Configuration: The experiment was based on GPT-4o-mini, with the Betti number fixed at 7 and mapping set to 3. The nine test datasets included GSM8K, SVAMP, MultiArith, ASDiv, AQUA (mathematical reasoning), MMLU (generalization), BBH, Date (multi-task QA), and CLUTRR (logical reasoning). The temperature parameter was adjusted from 0.0 to 1.0 in steps of 0.1, and the performance of SoT and CoT was compared. Five reasoning paths were generated in each experiment, and the average accuracy was recorded to reduce the impact of randomness.
Experimental Results:As shown in Figure 6, the accuracy of SoT across multiple datasets varies minimally with temperature (from 0.0 to 1.0), exhibiting stability with small amplitude. Figure 8indicates that the accuracy of CoT fluctuates significantly within the same temperature range, with the amplitude increasing as the temperature rises, especially at high temperatures (close to 1.0), where the fluctuations become more pronounced. The boxplot in Figure 7further compares the accuracy distributions of both methods across datasets: the box of SoT is relatively small, with few outliers, and its accuracy is concentrated with high consistency across datasets, even at high temperatures, without significant outliers. In contrast, the box of CoT is larger, with a more scattered distribution and more outliers, especially at high temperatures, where the reasoning consistency diminishes. Analysis suggests that the high temperature-induced diversity in CoT weakens logical coherence, while SoT maintains reasoning stability.
5. conclusion
In this paper, we introduce the SoT reasoning framework, integrating the MFR principle from algebraic geometry and homological algebra into the CoT paradigm, significantly enhancing LLM performance in complex reasoning tasks. SoT addresses traditional CoT limitations in high-dimensional, nonlinear, and logical scenarios by employing a structured decomposition strategy—utilizing Module, Freeness, Mapping, Exactness, Minimality, and Betti numbers—to create compact, logically consistent reasoning units.
While SoT has shown excellent reasoning capabilities with LLMs, further exploration is needed. We plan to extend the topological decomposition to multimodal reasoning, including inputs like images and tables, to verify its adaptability in cross-modal tasks. Additionally, incorporating iterative MFR concepts will enhance step-by-step problem refinement, optimizing reasoning paths dynamically and improving SoT’s efficiency and accuracy in multimodal and high-dimensional tasks.